OpenStudy (anonymous):

What is the next term of the following sequence? 12, 6, 3, ... 1 1/2?..

OpenStudy (anonymous):

it is in gp

OpenStudy (anonymous):

r = 6/12 = 1/2

OpenStudy (anonymous):

Every term is divided by 2 so: 3/2 = ???

OpenStudy (anonymous):

yes u r correct

OpenStudy (anonymous):

Yes you are correct...

OpenStudy (anonymous):

1 1/2 = 3/2